OpenStudy (anonymous):

Does anyone know how to integrate this? ∫2x^2-x/x^1/2..I need to solve this without using the substitution rule. any help would be greatly appreciated

OpenStudy (lgbasallote):

\[\int \frac{2x^2 - x}{x^{1/2}}dx?\]

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

first step would be to separate the integral \[\int \frac{2x^2}{x^{1/2}}dx - \int \frac{x}{x^{1/2}}dx\] did you do this?

OpenStudy (anonymous):

yea but i dont know what to do after

OpenStudy (lgbasallote):

next you apply the rule of exponents \[\frac{a^m}{a^n} \implies a^{m-n}\] remember that rule?

OpenStudy (anonymous):

oh yeea

OpenStudy (lgbasallote):

so do you know how to do the rest now?

OpenStudy (lgbasallote):

or do you need more help?

OpenStudy (anonymous):

i think i got it now. thanks a lot :)

OpenStudy (lgbasallote):

\[\huge \color{purple}{\textbf{<tips hat>}}\]